A 20g bullet moving at 30m/s stops after penetrating 3 cm bone . Calculate the average force exerted by the bullet .

Grade:11

1 Answers

Digpal singh
7 years ago
Time= distance/velocity= 3/(100)*(30)= 1/1000s. Force= mass*acceleration= 20/1000 * 30/1/1000 therefore, force= 20/1000 * 30000= 600Newton.